OpenStudy (anonymous):

1/square root of 80z

OpenStudy (anonymous):

no not correct

OpenStudy (anonymous):

\[\frac{1}{\sqrt{80z}}\times \frac{\sqrt{80z}}{\sqrt{80z}}=\frac{\sqrt{80z}}{80z}\]\]